Two cells each having an emf of 1.5v and an internal resistance of 2 ohms are connected in parallel. Find the current in each case when the cells are connected to a 1 ohm’s resister
1
Expert's answer
2020-11-06T10:13:59-0500

Current from one cell

"\Epsilon=I \times R_{c}+I \times R"

"I=\frac{E}{R_{c}+R}=\frac{1.5}{2+1}=0.5A"

Current from second cell is the same. They are both similar.

Current in resistor (summ of currents from two cells):

"I_{r}=I_{1}+I_{2}=0.5+0.5=1A"